Consolidation at date of acquisition (purchase price greater than book value, acquisition journal entries, deferred tax liability)
Assume that the parent company acquires its subsidiary by exchanging 116,000 shares of its $1 par value Common Stock, with a market value on the acquisition date of $30 per share, for all of the outstanding voting shares of the investee. In its analysis of the investee company, the parent values all of the subsidiary's assets and liabilities at an amount equaling their book values except for PPE assets that are undervalued by $1 million, an unrecorded Customer List that the parent values at $200,000, and an unrecorded Brand Name asset valued at $500,000. In addition, the purchase and sale agreement provides for contingent consideration in the event that the subsidiary meets earnings targets. You feel that the fair value of that contingent consideration liability is $400,000 on the acquisition date. And, finally, assume that the tax basis of the PPE assets is equal to their book values, and your tax rate is 35%.

a. Prepare the journal entry that the parent makes to record the acquisition.

General Journal
Description Debit Credit
Equity investmentCustomer listBrand nameGoodwillContingent consideration liabilityCommon stockAPICRetained earnings
Equity investmentCustomer listBrand nameGoodwillContingent consideration liabilityCommon stockAPICRetained earnings
APIC

b. Given the following acquisition-date balance sheets for the parent and its subsidiary, prepare the consolidation spreadsheet.

Elimination Entries
Balance Sheet Parent Subsidiary Dr Cr Consolidated
Assets
Cash $783,300 $104,000
Accounts receivable 384,000 696,000
Inventory 582,000 894,000
Equity investment 3,480,000 [E]
[A]
Property, plant and equipment (PPE), net 14,899,600 1,654,000 [A]
Customer list [A]
Brand name [A]
Equity investmentCustomer listBrand nameGoodwillContingent consideration liabilityCommon stockAPICRetained earnings [A]
$20,128,900 $3,348,000
Liabilities and stockholders' equity
Accounts payable $188,100 $127,000
Accrued liabilities 220,800 221,000
Long-term liabilities 2,000,000 1,000,000 [A]
Common stock 680,000 200,000 [E]
APIC 5,200,000 250,000 [E]
Retained earnings 11,840,000 1,550,000 [E]
$20,128,900 $3,348,000
Answer #1
Answer to (a)
Date Account Title & Explanation Debit ($) Credit ($)
Equity Investment (116000 x $30) 3480000
Common Stock (116000 x $1) 116000
Paid in capital (116000 x $29) 3364000
(Acquisition of shares)
Answer to (b)
Elimination Entries
Balance Sheet Parent Subsidiary Dr Cr Consolidated
Assets
Cash $783,300 $104,000 887300
Accounts receivable 384,000 696,000 1080000
Inventory 582,000 894,000 1476000
Equity investment 3,480,000 2000000 [E]
1880000 [A]
Property, plant and equipment (PPE), net 14,899,600 1,654,000 [A] 1000000 17153600
Customer list [A] 200000 200000
Brand name [A] 500000 500000
Equity investmentCustomer listBrand nameGoodwillContingent consideration liabilityCommon stockAPICRetained earnings [A] 180000 180000
$20,128,900 $3,348,000 $21,476,900
Liabilities and stockholders' equity
Accounts payable $188,100 $127,000 315100
Accrued liabilities 220,800 221,000 441800
Long-term liabilities 2,000,000 1,000,000 3000000
Deferred Tax liability -595000 [A] -595000
Common stock 680,000 200,000 [E] 200000 680000
APIC 5,200,000 250,000 [E] 250000 5200000
Retained earnings 11,840,000 1,550,000 [E] 2145000 12435000
$20,128,900 $3,348,000 $2,595,000 ($5,595,000) $21,476,900
Notes:
1)Deferred Tax Liability = (1000000 + 200000 + 500000) x 35/100 = $595000.
2)Total Equity Investments = (116000 x $30) + $400000 = $3880000
Equity Investments = Total Equity investments - PPE - Customer List - Brand name - Goodwill
Equity Investment = 3880000 - 1000000 - 200000 - 500000 - 180000 = $2000000